What are the steps to solve a system?
Here's the example:
x-2y=-7
5y=5x+20
I Have tried canceling out the -2y and 5y or 5x and x but I didn't get the answer right.
I Need the answer and steps so i know how to do this problem on my own when i have a test.
Please :)

Answer by stanbon(75887)

x-2y=-7
5y=5x+20
------
Solving by elimination:
x - 2y = -7
x - y = -4
--------
Subtract and solve for "y":
y = 3
----
Now solve for "x":
x - 2y = -7
x -2(3) = -7
x-6 = -7
x = -1
-----
Solution: (-1,3)
=====================
Cheers,
Stan H.
=====================